# Many-Body Localization in the Age of Classical Computing

### v2

@@ X.6 Higher Dimensions
Studies of ergodicity breaking phenomena in more than one spatial dimension are quite rare by comparison with the abundant literature for the 1D problem. The first stimulus for the consideration of MBL problem in higher dimensions came with experiments [128] in which half of a 2D lattice was initially filled with bosonic particles. Then the particles were allowed to wander in the presence of disorder, filling the lattice almost uniformly for a small disorder. In contrast, for a sufficiently strong disorder, the bosons remained in the vicinity of their initial positions and thermalization of the system was not achieved during the course of the experiment. The full scale simulation of the problem is beyond reach of the present classical computers. A first approach [712] used the time-dependent Gutzwiller approximation to tackle the problem. In this approach the system’s wave function remains separable throughout the evolution time so no entanglement can build up. The results of the simulation gave, however, quite an accurate description of the experiment showing that its interpretation in terms of MBL may be questioned, and the results can be interpreted in terms of slow glassy dynamics. The attempt at the simulation of full many-body dynamics of this bosonic system was undertaken on $6\times 6$ lattice with tree tensor network states by [670]. The authors simulated the experiment tuning the parameters of the problem to match the experimental system and, despite following the evolution for 256 tunneling times, they did not observe a saturation of the imbalance.
@@ X.7 Ergodicity breaking without disorder
Recent years brought a rich variety of interacting systems that reveal non-ergodic dynamics even in the absence of the disorder. The generic feature linking these models is the existence of additional approximate constants of motion that prevent thermalization at accessible time scales. The examples include breakdown of ergodicity in:
  • A. Lattice gauge theory models where typically the Gauss law, or other non-local conservation laws, provide restrictions for the dynamics [87, 233, 114];
  • B. Strong interactions that slow the dynamics and may lead to the Hilbert space fragmentation [546, 311];
  • C. Tilted lattices where emerging approximate conservation of the dipole moment leads to the so-called Stark MBL [446, 567].
The short-range models with the dynamical lattice gauge fields [233, 114], after the elimination of the fields with the Gauss law, become effectively long-ranged. This leads to MBL-like phenomenology. Understanding the possibilities of a breakdown of ergodicity in such models in the asymptotic limit $L,t\to\infty$ is an interesting direction for further research.
@@
The tilted lattice systems, i.e., the systems in which, instead of the on-site disorder, a linear potential $F\sum_{j}j\hat{n}_{i}$ is added to the Hamiltonian, provide yet another platform to study ergodicity breaking phenomena. The tile in optical lattices may be realized with an additional electric or space dependent magnetic field via Zeeman effect. Such a linear tilt of a 1D lattice leads, in the non-interacting limit, to the Wannier-Stark localization. Similarly, in the presence of interactions, 1D lattice models with tilted potential show signatures of non-ergodicity due to the Stark MBL [446, 567, 643]. For strong lattice tilts, the observed ergodicity breaking may be linked with the fragmentation of the Hilbert space due to the emerging dipole moment $\mathcal{D}=\sum_{j}j\hat{n}_{i}$ conservation, which recently stimulated a vivid interest in the subject [546, 518, 718, 112, 717]. Perturbing the titled potential with a weak disorder [446] or adding a small (e.g. quadratic) modification of the external potential [567, 643] results in the phenomenology similar to the one observed in the MBL regime of disordered spin chains, e.g., logarithmic entanglement entropy growth from initial separable states. Nevertheless, for a pure linear tilt a different behavior is observed [718, 716]. This may be linked to the emerging global dipole moment conservation in the system [172, 442]. Finally, we note that arguments for finite spin transport in tilted lattices were presented in [330], relying on a combination of mirror and spin-flip symmetries of such potential, as well as on the assumption of the absence of degeneracies in the spectrum.
@@ XI MBL in experiments
In this Section we give a brief overview of the experiments related to MBL. We highlight investigated system sizes and evolution times to illustrate the length and times scales relevant for the present-day experiments.
@@ XI.1 Early developments
Early years of this millennium were rich in efforts to experimentally verify Anderson localization, when it was understood that the cold atomic settings provide a new and promising medium for explorations of disordered many-body systems [151]. The early attempts to observe the Anderson localization were spoiled by weak interactions [133, 564, 210, 565]. Finally, diluting the cold atomic gas to make it non-interacting finally led to an experimental verification of the Anderson localization [75, 524].
Observation of MBL requires a preparation of the appropriate highly out-of equilibrium state with the finite energy density. In other words, the experiment should not be limited to low energy states as, e.g., in the first notable attempt to study the influence of strong disorder on the interacting bosons trapped in a 3D optical lattice [705]. Using the speckle disorder potential the authors were able to study the influence of the disorder on superfluid and Mott insulator phases of interacting bosons, as well as the possible emergence of the Bose glass phase. The latter study was extended later in [474].
A subsequent experiment [334] realizing disordered 3D Fermi-Hubbard model probed transport properties of the system. An external force was applied to displace the atoms from the centre of harmonic trapping potential and the velocity $v$ of their center-of-mass was measured. At weak disorder strength $W$, the velocity $v$ was manifestly non-vanishing suggesting presence of transport in the model. In contrast, large disorder strength $W$ lead to practical vanishing of $v$ suggesting absence of the transport. These observations are consistent with general expectations about the slowing down of the dynamics and the emergence of MBL regime with increasing the disorder strength $W$. However, their relation to the physics of thermalization and ergodicity breaking in quantum many-body system is not fully clear.
@@ XI.2 MBL regime in ultracold atomic systems
The breakthrough in studies of thermalization and MBL came with the experiment [561], in which the 1D Fermi-Hubbard model (88) was realized by a 3D optical lattice with deep confining potential in two spatial dimensions leading to an array of decoupled 1D tubes. The disorder was added via a second weak laser, incommensurate with the one creating lattices, resulting in a quasiperiodic potential such as the one in Eq. (91). A preparation of the initial highly inhomogeneous state at high energy density with every second lattice site occupied by a fermion and every second site empty was followed by monitoring of the time evolution of the system. Measuring the density autocorrelation function (a generalization of $C(t)$ in Eq. (46) for spinful fermions), and varying the interaction strength $U$ and the amplitude of quasiperiodic potential $W$, the experiment [561] demonstrated a crossover between the ergodic (thermalizing) behavior, for which vanishing of $C(t)$ is observed, and the MBL regime, in which $C(t)$ is significantly non-zero even at the longest probed time scale.
@@
A subsequent experiment [399] extended the available time scale to $t_{\mathrm{max}}=100$, which allowed for observation of the slow, power-law like, decays of the density autocorrelation function $C(t)$ in the vicinity of the MBL regime. The absence of the Griffiths regions in a quasiperiodic potential considered by [399] indicates that a different type of mechanism may be responsible for the decay of $C(t)$, see Sec. VII.3.2. Introduction of the coupling between the 1D tubes significantly enhances the tendencies towards thermalization, increasing the disorder strengths needed to observe the MBL regime [83, 81]. Experiments in this line also considered the ETH-MBL crossover in the periodically driven MBL systems [82] and showed how the MBL regime is destroyed in the presence of an engineered dissipation [398].
The 1D Bose-Hubbard model (89) with quasiperiodic potential was realized in a line of experiments with a quite small number of ultracold bosons in optical lattices [397, 523, 361]. The exquisite control of the setup enabled experimental studies of the slow growth of entanglement entropy from an initial product state in the MBL regime [397] (see Sec. VII.2), as well as probing 2-point correlation functions at the ETH-MBL crossover [523]. The single-site resolution enabled observation of signatures of quantum avalanches by planting thermal inclusions [361], as described in Sec. IX.1.3. The typical lattice size considered in this experimental realization of the Bose-Hubbard Hamiltonian is $L=12$ and the probed time scales are of the order of $t_{\mathrm{max}}=100$ tunneling times. A separate experiment realized disordered Bose-Hubbard model [128] on a 2D lattice comprising in total $L\approx 500$ sites. Studies of time evolution of the density autocorrelation function $C(t)$ for initially prepared out-of-equilibrium state, in which bosons were occupying one half of the lattice, have demonstrated a crossover between thermalization at weak disorder strengths and MBL regime at strong disorder. In the latter regime, the bosons remained in the vicinity of their initial positions even after $t_{\mathrm{max}}=300$ tunneling times.
For completeness, we mention realization of a titled Fermi-Hubbard model, which hosts non-ergodic regimes due to Hilbert space fragmentation [554, 333]. Here up to $t_{\mathrm{max}}=700$ tunneling times have been reached. Experimental realization of quasicrystals [574] in systems of ultracold bosons in 2D optical lattices [553, 720] opens up a new testing ground for the MBL regime in 2D, especially interesting due to the absence of rare-region effects in these systems.
@@ XI.3 MBL regime in other platforms
Quantum Ising model (87) with additional long-range couplings was realized in Yb-ion quantum simulator with programmable random disorder [614]. The experiment with $L=10$ spins has shown a crossover between thermalization at weak disorder and the presence of MBL regime at large disorder in which the density autocorrelation functions do not decay at the longest considered time scale $t_{\mathrm{max}}=10$. Assuming the relevance of the quantum avalanche mechanism the observed MBL regime should eventually become ergodic with increasing time and length scales, see Sec. X.5.
Disordered 1D Bose-Hubbard was also implemented on an $L=9$ superconducting qubit system [529], and the ETH-MBL crossover was observed experimentally in the spectral properties of the system. In similar spirit, the experiment on up to $L=21$ superconducting qubits [126] where the tomographic reconstruction of single and double qubit density matrices helped to quantify temporal and spatial aspects of the entanglement growth in the MBL regime, explored the time scales up to $t_{\mathrm{max}}\approx 400$ tunneling times. An experiment monitoring the dynamics of $L=19$ superconducting qubits over about $t_{\mathrm{max}}\approx 100$ tunneling times [253] demonstrated that the disorder strength at which the ETH-MBL crossover occurs is dependent on the energy of the initial state [110]. Fock-space perspective on the dynamics in ergodic and MBL regimes, see Sec. VI.2.1, was investigated experimentally in a system of $L=24$ superconducting qubits [714].
@@ XI.4 Perspective on MBL experiments
The numerical results reviewed in this work typically concern time scales and system sizes similar to or larger than those accessible in present-day experiments. The notable exception is provided by cold atomic settings in which the system sizes significantly exceed the capabilities of exact numerical methods. Nevertheless, even in those cases, the finite coherence times allow us to reproduce the significant features of the experiments, at least at larger disorder strengths with tensor network approaches. The numerical results, while indicating the possible benefits of increasing the coherence times and experimentally probed system sizes, at least partially discourage such an approach. Let us consider a gedanken experiment in which the unitary time evolution of a disordered many-body system can be probed to $t_{\mathrm{max}}=10^{4}$ tunneling times in a lattice comprising $L=1000$ qubits. Can we expect the emergence of any new trend in the density autocorrelation function $C(t)$ in such an experiment? Sec. VII suggests that the answer is no. Fixing disorder strength $W$ and studying the time evolution of $C(t)$, we would likely find that $C(t)$ decays to smaller values at the longest evolution time than in an experiment with a smaller system and shorter probed times. Does it mean that no further progress is possible? Again, we believe that progress is indeed possible. Investigating microscopic mechanisms responsible for the emergence and breakdown of ergodicity, including, but not limited to, the scenarios considered in Sec. IX, is experimentally feasible and may lead to more conclusive answers about the MBL phase.
@@ XII.1 Summary
We close this review by revisiting the open questions outlined in Sec. IV.1. We comment on these problems in light of the numerical results for disordered interacting quantum many-body systems reviewed above.
1. Definition of MBL. The MBL regime, in which interacting many-body systems appear to avoid thermalization, is clearly observed for sufficiently large disorder strength $W$ in the systems of size $L$ studied numerically and experimentally. The breakdown of thermalization of local observables in the asymptotic limit $L,t\to\infty$, Eq. (6), is a natural, from a statistical physics perspective, way to define the MBL phase.
@@
One option for an alternative definition of the MBL phase involves constraining the probed times $t$ by a certain threshold. For instance, one could consider only times scaling no faster than polynomially with the system size $L$. Restoration of the ergodicity at longer times, e.g., scaling exponentially with $L$, would be irrelevant from the point of view of such a definition, and systems that fail to thermalize at times scaling polynomially with $L$ would belong to some form of non-ergodicity, e.g., to the MBL phase. However, the abrupt slow-down of the dynamics of disordered systems makes such a definition of MBL not much more practical than the definition involving the $L,t\to\infty$ limit. Indeed, the decay of local correlators $C(t)$ at disorder $W=8$ in the XXZ spin chain is well approximated by a power-law behavior $C(t)\propto t^{-\beta}$, where $\beta\approx 10^{-3}$. If we allow the time to be a polynomial of $L$ of order higher than $\frac{1}{\beta}$, the XXZ chain at $W=8$ may still be ergodic. However, to verify this, one would have to wait for an enormous amount of time, which is, again, irrelevant from the experimental perspective. Therefore, the question of proposing a definition of the MBL phase that would be both practically relevant and satisfactory from the theoretical standpoint remains a largely open problem.
2. Understanding of the MBL phase. Assuming that the MBL phase is defined by the failure of thermalization in the asymptotic limit $L,t\to\infty$, presently available numerical (and experimental) results do not provide a clear-cut understanding of the MBL phase. A broad consensus among mathematicians and physicists regarding rigorous results about the status of the MBL phase defined in that manner is yet to be achieved.
@@
4. Mechanism of the transition. The quantum avalanche scenario, which describes thermalization of the MBL region due to rare ergodic inclusions emerging in regions of anomalously weak disorder, is the prevalent proposition for a microscopic mechanism of the MBL transition. As we argued in Sec. IX, the implications of the avalanche mechanism can be studied numerically and experimentally, particularly by engineering conditions for the propagation of avalanches by planting thermal inclusions. Determination of the relevance of the avalanche mechanism for the thermalization in disordered many-body systems and exploration of its falsifiable consequences are important directions for further research. This includes comprehending how to model the effects of the rare thermal inclusions in a simplified manner, and what aspects of the dynamics of isolated disordered many-body systems may be captured by the open system approach.
Many-body resonances arise when eigenstates of disordered many-body systems are superpositions of a few product states that differ locally at extensively many lattice sites. The MBL transition may be associated with the appearance of system-wide many-body resonances that corroborate a breakdown of locality in the disordered spin chain. Establishing links between the quantities that probe rare instances of the many-body resonances, described in Sec. IX.2.3 and  IX.2.4 and the standard ergodicity breaking indicators involving averages of quantities derived from spectra of many-body systems, discussed in Sec. VI, is an essential direction for further explorations.
@@
6. Models of MBL and beyond. Disordered XXZ spin-1/2 chain and its close cousin, the disordered $J_{1}$-$J_{2}$ model, serve as paradigmatic models for studies of MBL. The majority of the numerical results reviewed here concern these models. Many disordered 1D many-body systems, mentioned in Sec. X, share similar dynamical features, including finite size drifts at the ETH-MBL crossover and the regimes of persistent slow dynamics at strong disorder. Despite the similarities, details of the numerical results vary from model to model. For instance, the MBL phenomenology in the disordered spin-1/2 chains appears more stable at available system sizes in Floquet systems than in Hamiltonian systems. In contrast, when spin-1/2 chains’ dynamics are subject to specific local constraints, numerical results indicate much stronger tendencies towards ergodicity, suggesting a lack of the MBL phase in such systems. These disparities indicate that the status of MBL may depend on the microscopic details of particular many-body Hamiltonians. Identifying and classifying microscopic features that enhance the ergodic phase or lead to more robust premises for the occurrence of the MBL phase is an important avenue for further studies. Such efforts can be facilitated by studies of interacting models in which the ergodicity-breaking transition is more accessible numerically, such as the quantum sun model introduced within the avalanche theory.

### v3

@@ X.6 Higher Dimensions
Studies of ergodicity breaking phenomena in more than one spatial dimension are quite rare by comparison with the abundant literature for the 1D problem. The first stimulus for the consideration of MBL problem in higher dimensions came with experiments [133], in which half of a 2D lattice was initially filled with bosonic particles. Then the particles were allowed to wander in the presence of disorder, filling the lattice almost uniformly for a small disorder. In contrast, for a sufficiently strong disorder, the bosons remained in the vicinity of their initial positions and thermalization of the system was not achieved during the course of the experiment. The full scale simulation of the problem is beyond reach of the present classical computers. A first approach [768] used the time-dependent Gutzwiller approximation to tackle the problem. In this approach the system’s wave function remains separable throughout the evolution time so no entanglement can build up. The results of the simulation gave, however, quite an accurate description of the experiment showing that its interpretation in terms of MBL may be questioned, and the results can be interpreted in terms of slow glassy dynamics. The attempt at the simulation of full many-body dynamics of this bosonic system was undertaken on a $6\times 6$ lattice with tree tensor network states by [724]. The authors simulated the experiment tuning the parameters of the problem to match the experimental system and, despite following the evolution for 256 tunneling times, they did not observe a saturation of the imbalance.
@@ X.7 Ergodicity breaking without disorder
Recent years brought a rich variety of interacting non-integrable systems that reveal non-ergodic dynamics even in the absence of the disorder. The generic feature linking these models is the existence of additional approximate constants of motion that prevent thermalization at accessible time scales. The examples include breakdown of ergodicity in:
  • A. Lattice gauge theory models where typically the Gauss law, or other local conservation laws, provide restrictions for the dynamics [90, 249, 118];
  • B. Strong interactions that slow down the dynamics and may lead to the Hilbert space fragmentation [592, 337];
  • C. Tilted lattices where the emerging approximate conservation of the dipole moment leads to the so-called Stark MBL [484, 614].
The short-range models with the dynamical lattice gauge fields [249, 118], after the elimination of the fields with the Gauss law, become effectively long-ranged. This leads to the MBL-like phenomenology. Understanding the possibilities of a breakdown of ergodicity in such models in the asymptotic limit $L,t\to\infty$ is an interesting direction for further research.
@@
The tilted lattice systems, i.e., the systems in which, instead of the on-site disorder, a linear potential $F\sum_{j}j\hat{n}_{i}$ is added to the Hamiltonian, provide yet another platform to study ergodicity breaking phenomena. The tile in optical lattices may be realized with an additional electric or space dependent magnetic field via Zeeman effect. Such a linear tilt of a 1D lattice leads, in the non-interacting limit, to the Wannier-Stark localization. Similarly, in the presence of interactions, 1D lattice models with tilted potential show signatures of non-ergodicity due to the Stark MBL [484, 614, 695]. For strong lattice tilts, the observed ergodicity breaking may be linked to the fragmentation of the Hilbert space due to the emerging conservation of a dipole moment, $\mathcal{D}=\sum_{j}j\hat{n}_{i}$, which recently stimulated a vivid interest in the subject [592, 559, 774, 116, 773, 480]. Perturbing the titled potential with a weak disorder [484] or adding a small (e.g. quadratic) modification of the external potential [614, 695] results in the phenomenology similar to the one observed in the MBL regime of disordered spin chains, e.g., logarithmic entanglement entropy growth from initial separable states. Nevertheless, for a pure linear tilt a different behavior is observed [774, 772]. This may be linked to the emerging global dipole moment conservation in the system [181, 480]. Finally, we note that arguments for finite spin transport in tilted lattices were presented in [358], relying on a combination of mirror and spin-flip symmetries of such potential, as well as on the assumption of the absence of degeneracies in the spectrum.
Experimental investigations of Rydberg atom arrays [64, 97] lead to the discovery of a form of weak-ergodicity breaking, defined via a strong dependence of the thermalization process on the system’s initial state. For instance, the PXP model (90), for $\alpha=1$ and $h_{i}=0$, initialized in the Néel state $|\ket{\uparrow\downarrow\uparrow\downarrow\ldots}$, features long-lived oscillations of local observables and revivals of wave function, while, for period-4 density wave state $|\ket{\uparrow\downarrow\downarrow\downarrow\uparrow\ldots}$, the thermalization occurs quickly [722]. This dynamical behavior of the PXP model is due to quantum many-body scar states [723]. The quantum scar states form a set of equally spaced in energy eigenstates of $\hat{H}_{\mathrm{PXP}}$ characterized by low entanglement and atypically large overlaps with the Néel state. The set of scar states is embedded in the spectrum of ergodic eigenstates of $\hat{H}_{\mathrm{PXP}}$. Ideas to employ quantum many-body scars as a mechanism for maintaining coherence and the desire for a more complete understanding of dynamics of non-integrable many-body systems fueled the intensive investigations of weak-ergodicity breaking phenomena, as reviewed in [622, 469, 120]. Non-thermal eigenstates can be found in multiple non-integrable quantum models, including Fermi-Hubbard model [313, 471], Bose-Hubbard model [309], spin-1 systems [600, 123], spin-1/2 model with emergent kinetic constraints [312], and even in disordered spin chains [634]. The emergence of a decoupled subspace spanned by non-thermal eigenstates in an otherwise thermal spectrum of the system [635] is a common trait of the models of weak-ergodicity breaking.
@@ XI MBL in experiments
In this Section we give a brief overview of the experiments related to MBL. We highlight the investigated system sizes and evolution times to illustrate the length and times scales relevant for the present-day experiments.
@@ XI.1 Early developments
Early years of this millennium were rich in efforts to experimentally verify Anderson localization, when it was understood that the cold atomic settings provide a new and promising medium for explorations of disordered many-body systems [157]. The early attempts to observe the Anderson localization were spoiled by weak interactions [138, 611, 224, 612]. Finally, diluting the cold atomic gas to make it non-interacting, led to an experimental verification of the Anderson localization [78, 569].
Observation of MBL requires a preparation of the appropriate highly out-of equilibrium state with the finite energy density. In other words, the experiment should not be limited to low energy states as, e.g., in the first notable attempt to study the influence of strong disorder on the interacting bosons trapped in a 3D optical lattice [761]. Using the speckle disorder potential, the authors were able to study the influence of the disorder on superfluid and Mott insulator phases of interacting bosons, as well as the possible emergence of the Bose glass phase. The latter study was extended later in [512].
A subsequent experiment, [362] realized the disordered 3D Fermi-Hubbard model and probed transport properties of the system. An external force was applied to displace the atoms from the centre of harmonic trapping potential and the velocity $v$ of their center-of-mass was measured. At weak disorder strength $W$, the velocity $v$ was manifestly non-vanishing, suggesting the presence of transport in the model. In contrast, large disorder strength $W$ lead to practical vanishing of $v$ suggesting absence of the transport. These observations are consistent with general expectations about the slowing down of the dynamics and the emergence of MBL regime with increasing the disorder strength $W$. However, their relation to the physics of thermalization and ergodicity breaking in quantum many-body system is not fully clear.
@@ XI.2 MBL regime in ultracold atomic systems
The breakthrough in studies of thermalization and MBL came with the experiment [608], in which the 1D Fermi-Hubbard model (88) was realized by a 3D optical lattice with deep confining potential in two spatial dimensions leading to an array of decoupled 1D tubes. The disorder was added via a second weak laser, incommensurate with the one creating lattices, resulting in a quasiperiodic potential such as the one in Eq. (91). A preparation of the initial highly inhomogeneous state at high energy density, with every second lattice site occupied by a fermion and every second site empty, was followed by monitoring of the time evolution of the system. Measuring the density autocorrelation function (a generalization of $C(t)$ in Eq. (46) for spinful fermions), and varying the interaction strength $U$ and the amplitude of quasiperiodic potential $W$, the experiment [608] demonstrated a crossover between the ergodic (thermalizing) behavior, for which vanishing of $C(t)$ is observed, and the MBL regime, in which $C(t)$ is significantly non-zero even at the longest probed time scale.
@@
A subsequent experiment [435] extended the available time scale to $t_{\mathrm{max}}=100$, which allowed for observation of the slow, power-law like, decays of the density autocorrelation function $C(t)$ in the vicinity of the MBL regime. The absence of the Griffiths regions in a quasiperiodic potential considered by [435] indicates that a different type of mechanism may be responsible for the decay of $C(t)$, see Sec. VII.3.2. Introduction of the coupling between the 1D tubes significantly enhances the tendencies towards thermalization, increasing the disorder strengths needed to observe the MBL regime [86, 84]. Experiments in this line also considered the ETH-MBL crossover in the periodically driven MBL systems [85] and showed how the MBL regime is destroyed in the presence of an engineered dissipation [434], following a theoretical analysis by [398, 217].
The 1D Bose-Hubbard model (89) with quasiperiodic potential was realized in a line of experiments with a quite small number of ultracold bosons in optical lattices [433, 568, 393]. The exquisite control of the setup enabled experimental studies of the slow growth of entanglement entropy from an initial product state in the MBL regime [433] (see Sec. VII.2), as well as probing 2-point correlation functions at the ETH-MBL crossover [568]. The single-site resolution enabled observation of signatures of quantum avalanches by planting thermal inclusions [393], as described in Sec. IX.1.3. The typical lattice size considered in this experimental realization of the Bose-Hubbard Hamiltonian is $L=12$ and the probed time scales are of the order of $t_{\mathrm{max}}=100$ tunneling times. A separate experiment realized disordered Bose-Hubbard model [133] on a 2D lattice comprising in total $L\approx 500$ sites. Studies of time evolution of the density autocorrelation function $C(t)$ for the initially prepared out-of-equilibrium state, in which bosons were occupying one half of the lattice, demonstrated a crossover between thermalization at weak disorder strengths and MBL regime at strong disorder. In the latter regime, the bosons remained in the vicinity of their initial positions even after $t_{\mathrm{max}}=300$ tunneling times.
For completeness, we mention realization of a titled Fermi-Hubbard model, which hosts non-ergodic regimes due to Hilbert space fragmentation [601, 361]. Here, up to $t_{\mathrm{max}}=700$ tunneling times have been reached. Experimental realization of quasicrystals [621] in systems of ultracold bosons in 2D optical lattices [599, 776] opens up a new testing ground for the MBL regime in two dimensions, especially interesting due to the absence of rare-region effects in these systems.
@@ XI.3 MBL regime in other platforms
Quantum Ising model (87) with additional long-range couplings was realized in Yb-ion quantum simulator with programmable random disorder [664]. The experiment with $L=10$ spins has shown a crossover between thermalization at weak disorder and the presence of MBL regime at large disorder in which the density autocorrelation functions do not decay at the longest considered time scale $t_{\mathrm{max}}=10$. Assuming the relevance of the quantum avalanche mechanism, the observed MBL regime should eventually become ergodic with increasing time and length scales, see Sec. X.5.
Disordered 1D Bose-Hubbard model was also implemented on an $L=9$ superconducting qubit system [575], and the ETH-MBL crossover was observed experimentally in the spectral properties of the system. In similar spirit, the experiment on up to $L=21$ superconducting qubits [131], where the tomographic reconstruction of single and double qubit density matrices helped to quantify temporal and spatial aspects of the entanglement growth in the MBL regime, explored the time scales up to $t_{\mathrm{max}}\approx 400$ tunneling times. An experiment monitoring the dynamics of $L=19$ superconducting qubits over about $t_{\mathrm{max}}\approx 100$ tunneling times [269] demonstrated that the disorder strength at which the ETH-MBL crossover occurs is dependent on the energy of the initial state [114]. Fock-space perspective on the dynamics in ergodic and MBL regimes, see Sec. VI.2.1, was investigated experimentally in a system of $L=24$ superconducting qubits [770].
@@ XI.4 Perspective on MBL experiments
The numerical results reviewed in this work typically concern time scales and system sizes similar to or larger than those accessible in present-day experiments. The notable exception is provided by cold atomic settings in which the system sizes significantly exceed the capabilities of exact numerical methods. Nevertheless, even in those cases, the finite coherence times allow us to reproduce the significant features of the experiments, at least at larger disorder strengths, with tensor network approaches. The numerical results, while indicating the possible benefits of increasing the coherence times and experimentally probed system sizes, at least partially discourage such an approach. Let us consider a gedanken experiment in which the unitary time evolution of a disordered many-body system can be probed to $t_{\mathrm{max}}=10^{4}$ tunneling times in a lattice comprising $L=1000$ qubits. Can we expect the emergence of any new trend in the density autocorrelation function $C(t)$ in such an experiment? Sec. VII suggests that the answer is no. Fixing disorder strength $W$ and studying the time evolution of $C(t)$, we would likely find that $C(t)$ decays to smaller values at the longest evolution time than in an experiment with a smaller system and shorter probed times. Does it mean that no further progress is possible? Again, we believe that progress is indeed possible. Investigating microscopic mechanisms responsible for the emergence and breakdown of ergodicity, including, but not limited to, the scenarios considered in Sec. IX, is experimentally feasible and may lead to more conclusive answers about the MBL phase.
@@ XII.1 Summary
We close this review by revisiting the open questions outlined in Sec. IV.1. We comment on these problems in light of the numerical results for the disordered interacting quantum many-body systems reviewed above.
1. Definition of MBL. The MBL regime, in which interacting many-body systems appear to avoid thermalization, is clearly observed for sufficiently large disorder strength $W$ in the systems of size $L$ studied numerically and experimentally. The breakdown of thermalization of local observables in the asymptotic limit $L,t\to\infty$, Eq. (6), is a natural, from the statistical physics perspective, way to define the MBL phase.
@@
One option for an alternative definition of the MBL phase involves constraining the probed times $t$ by a certain threshold. For instance, one could consider only times scaling no faster than polynomially with the system size $L$. Restoration of the ergodicity at longer times, e.g., scaling exponentially with $L$, would be irrelevant from the point of view of such a definition, and systems that fail to thermalize at times scaling polynomially with $L$ would belong to some form of non-ergodicity, e.g., to the MBL phase. However, the abrupt slow-down of the dynamics of disordered systems makes such a definition of MBL not much more practical than the definition involving the $L,t\to\infty$ limit. Indeed, the decay of local correlators $C(t)$ at disorder $W=8$ in the disordered XXZ spin chain is well approximated by a power-law behavior, $C(t)\propto t^{-\beta}$, where $\beta\approx 10^{-3}$. If we allow the time to be a polynomial of $L$ of order higher than $\frac{1}{\beta}$, the disordered XXZ chain at $W=8$ may still be ergodic. However, to verify this, one would have to wait for an enormous amount of time, which is, again, irrelevant from the experimental perspective. Therefore, the question of proposing a definition of the MBL phase that would be both practically relevant and satisfactory from the theoretical standpoint remains a largely open problem.
2. Understanding of the MBL phase. Assuming that the MBL phase is defined by the failure of thermalization in the asymptotic limit $L,t\to\infty$, presently available numerical (and experimental) results do not provide a clear-cut understanding of the MBL phase. A broad consensus among mathematicians and physicists regarding rigorous results about the status of the MBL phase defined in that manner is yet to be achieved. The multi-scale Jacobi method considered in [317, 573] or other rigorous mathematical framework will be needed to finally resolve this question.
@@
4. Mechanism of the transition. The quantum avalanche scenario, which describes thermalization of the MBL region due to rare ergodic inclusions emerging in the regions of anomalously weak disorder, is the prevalent proposition for a microscopic mechanism of the MBL transition. As we argued in Sec. IX, the implications of the avalanche mechanism can be studied numerically and experimentally, particularly by engineering conditions for the propagation of avalanches by planting thermal inclusions. Determination of the relevance of the avalanche mechanism for the thermalization in disordered many-body systems and exploration of its falsifiable consequences are important directions for further research. This includes comprehending how to model the effects of the rare thermal inclusions in a simplified manner, and what aspects of the dynamics of isolated disordered many-body systems may be captured by the open system approach.
Many-body resonances arise when eigenstates of disordered many-body systems are superpositions of a few product states that differ locally at extensively many lattice sites. The MBL transition may be associated with the appearance of system-wide many-body resonances that corroborate a breakdown of locality in the disordered spin chain. Establishing links between the quantities that probe rare instances of the many-body resonances, described in Sec. IX.2.3 and  IX.2.4, and the standard ergodicity breaking indicators involving averages of quantities derived from spectra of many-body systems, discussed in Sec. VI, is an essential direction for further explorations.
@@
6. Models of MBL and beyond. Disordered XXZ spin-1/2 chain and its close cousin, the disordered $J_{1}$-$J_{2}$ model, serve as paradigmatic models for studies of MBL. The majority of the numerical results reviewed here concern these models. Many disordered 1D many-body systems, mentioned in Sec. X, share similar dynamical features, including the finite size drifts at the ETH-MBL crossover and the regimes of persistent slow dynamics at strong disorder. Despite the similarities, details of the numerical results vary from model to model. For instance, the MBL phenomenology in the disordered spin-1/2 chains appears more stable at available system sizes in Floquet systems than in Hamiltonian systems. In contrast, when spin-1/2 chains’ dynamics are subject to specific local constraints, numerical results indicate much stronger tendencies towards ergodicity, suggesting a lack of the MBL phase in such systems. These disparities indicate that the status of MBL may depend on the microscopic details of particular many-body Hamiltonians. Identifying and classifying microscopic features that enhance the ergodic phase or lead to more robust premises for the occurrence of the MBL phase is an important avenue for further studies. Such efforts can be facilitated by studies of interacting models in which the ergodicity-breaking transition is more accessible numerically, such as the quantum sun model introduced within the avalanche theory and the ultrametric model defined on a Fock space.